supposed 100 refrigerator were manufactured during a day. ninety five operated perfectly and 5 had atleast one defect. a sample of four is selected at random. using the hypergeometric distribution formula, find the probability that 8 of the 10 will operate perfectly?

1
Expert's answer
2022-02-22T00:38:41-0500

Solution:

  • K is the number of successes in the population, K=95
  • k is the number of observed successes,k=8
  • N is the population size, N=100
  • n is the number of draws, n=10

"P(X=k)=\dfrac{^KC_k\times ^{N-K}C_{n-k}}{^NC_n}
\\ P(X=8)=\dfrac{^{95}C_8\times ^{100-95}C_{10-8}}{^{100}C_{10}}
\\=\dfrac{^{95}C_8\times ^{5}C_{2}}{^{100}C_{10}}
\\=0.0702"